OpenStudy (anonymous):

How do I simplify i to the power of 23?

jimthompson5910 (jim_thompson5910):

Notice this pattern i^1 = i i^2 = -1 i^3 = -i i^4 = 1 i^5 = i i^6 = -1 i^7 = -i i^8 = 1 ... it keeps going and it repeats on every 4th number (starting from 1). That means i^1 = i^5 for instance or i^7 = i^3

jimthompson5910 (jim_thompson5910):

You don't have to keep writing out that pattern to get to i^23, you can use this rule below Step 1) Take the exponent over 'i' and divide it by 4 Step 2) The remainder you get will tell you what the answer is when you compare it to that pattern. ------------------------------------------------- So let's use these steps to compute i^23 Step 1) Divide the exponent 23 by 4: 23/4 = 5 remainder 3 Step 2) The remainder is 3. So that means i^23 = i^3 = -i. In short, i^23 = -i

jimthompson5910 (jim_thompson5910):

Another example: Let's say we want to find i^56 Step 1) Divide the exponent of 56 by 4 to get 56/4 = 14 remainder 0 Step 2) The remainder is 0, so i^56 = i^0. Because i^0 = 1, we know that i^56 = 1